# secReferences

### v1

@@ 5.1 Simulation under the random graph models
We first conduct a simulation study under the random graph models to evaluate the numerical performance of different fairness criteria. We consider a binary protected feature that indicates if one is white or non-white. Specifically, we fix 40 white incompatible pairs (80%) and 10 non-white incompatible pairs (20%). Among the 40 white pairs, 28 patients (70%) are lowly-sensitized, 8 patients (20%) are moderately-sensitized, and 4 patients (10%) are highly-sensitized; and among the 10 non-white pairs, 7 patients (70%) are lowly-sensitized, 2 patients (20%) are moderately-sensitized, and 1 patient (10%) is highly-sensitized. The PRA scores are set to be 0.9, 0.45 and 0.05 for highly-sensitized, moderately-sensitized, and lowly-sensitized patients, respectively. The distribution follows the analysis in Saidman et al. (2006).
@@
We aim to compare the three fairness criteria, group fairness, individual fairness, and our new fairness, with no fairness as a baseline result. We consider the two ways of parameter selection in Section 3.1 for our new fairness. For group fairness, we consider the two ways of parameter selection. The first way is to choose the largest possible $\alpha$ such that (2) is solvable, as discussed in Dickerson et al. (2014). This represents the strongest possible constraint that we must maximize the number of matched highly-sensitized patients. The second one chooses the largest possible $\alpha$ such that the overall utility is equal to that in (1), as discussed in Freedman et al. (2020). In other words, among all the exchange plans that maximize the objective function in (2), we consider the one that maximizes the number of matched highly-sensitized patients. For individual fairness, we choose the $L_{2}$ norm of the vector $(\delta_{v}-\overline{\delta})_{v\in V}$ and two candidate values $\{(0.15/N)^{1/2},(0.25/N)^{1/2}\}$ of $\beta$ in (3). That is, we require the sample variance of $(\delta_{v}-\overline{\delta})_{v\in V}$ to be no greater than $0.15$ and $0.25$, respectively; and the candidate value $(0.15/N)^{1/2}$ represents a stronger constraint and the candidate value $(0.25/N)^{1/2}$ represents a weaker constraint.
Figure 2 presents the average selection rates of different fairness criteria within each subgroup. From Figure 2, group fairness works to increase the selection rates of highly-sensitized patients, and individual fairness works to balance the selection rates of the six subgroups. Differently, our fairness balances the selection rates of the white and non-white patients within each sensitization stratum, instead of the selection rates of all six subgroups. Moreover, our fairness achieves the lowest average absolute difference in selection rates between the two race groups. For the price of fairness, although our fairness (strong) has the lowest average utility, the relative utility loss (28.7-27.91)/28.7=2.8% is relatively low.
@@ 5.2 UNOS data analysis
Figure 3 reports the average selection rates within each subgroup. From Figure 3, group fairness, individual fairness, and our new fairness approach all enhance equitable access to transplant opportunities according to their specific fairness criteria. However, compared to the idealized random graph model discussed in Section 5.1, the price of implementing our new fairness criteria is significantly higher, approximately 8.4%. This discrepancy arises because the PRA scores in the UNOS dataset can vary continuously between 0 and 1, whereas in the previous simulation study in Section 5.1, PRA scores were limited to three discrete values: 0.9, 0.45, and 0.05. Consequently, within each sensitization level, the probability of compatibility with a random donor can differ substantially, increasing the difficulty of balancing selection rates across subgroups. This indicates a potential limitation of our fairness criteria, suggesting the need for a more precise discretization of PRA scores.
@@ 5.3 Experiment on selection probability prediction
We next conduct a numerical experiment to evaluate the prediction accuracy of selection probability. We assume there is an underlying population consisting of 80% white and 20% non-white incompatible pairs, following the same blood type distributions and sensitization distributions in Section 5.1. We assume the historical data $\{\tilde{v}_{1},...,\tilde{v}_{N_{0}}\}$, current data $\{v_{1},...,v_{N_{1}}\}$, and future data are independently sampled from the population. We fix $N_{0}=200$, $N=100$, $B=1000$, and vary $N_{1}$ in $\{20,40,60,80\}$. For illustration purposes, we focus on our new fairness.
Table 1 presents the numerical results of the selection probability prediction. As $L$ increases, the prediction accuracy improves, evidenced by a lower mean squared error (MSE) and a coverage rate of the prediction interval approaching 95%. This improvement occurs because the size of the unobserved future data, $N-N_{1}$, decreases, which makes the prediction easier. Although these results indicate the method’s validity, there is a bias in the prediction interval due to the finite size of $N_{0}$. When $N_{0}$ is large, the set $\{\tilde{v}_{1},...,\tilde{v}_{N_{0}}\}$ closely approximates the underlying population distribution; while when $N_{0}$ is small, the distribution of $\{\tilde{v}_{1},...,\tilde{v}_{N_{0}}\}$ may differ from the underlying population distribution.
@@
MSE 0.087 0.076 0.073 0.059
Coverage 0.986 0.983 0.980 0.981
Width 0.589 0.566 0.541 0.522
Table 1: Accuracy of the predicted selection probability of $v_{1},...,v_{N_{1}}$. The results are averaged over 50 data replications.
@@ Discussion
Throughout this paper, we assume that the protected feature $A$ is binary. While it is possible to extend $A$ to a general categorical variable and derive results analogous to Proposition 1, such an extension would alter the upper bound of the number of exchange plans in $\mathcal{F}$ with nonzero selection probability, significantly increasing the computational complexity of the algorithm. We leave it for future research to explore more efficient algorithms to accommodate these potential extensions.
@@ Appendix A Proofs
For arbitrary $x_{1},...,x_{n}\in\mathbb{R}^{p}$, any boundary point of the convex hull of the point set $\{x_{1},...,x_{n}\}$ can be represented as $\sum_{i=1}^{n}t_{i}x_{i}$, satisfying that $t_{1},...,t_{n}\geq 0$, $\sum_{i=1}^{n}t_{i}=1$, and the number of nonzero elements in $t_{1},...,t_{n}$ is no greater than $p$.
@@
When $n\leq p$, the result is trial. We only consider the case when $n>p$ in the following proof.
For any point $y\in\mathbb{R}^{p}$ in the convex hull of the point set $\{x_{1},...,x_{n}\}$, we can write $y=\sum_{i=1}^{n}t_{i}x_{i}$, where $t_{1},...,t_{n}\geq 0$ and $\sum_{i=1}^{n}t_{i}=1$. There are different possible choices of $t_{1},...,t_{n}$ satisfying $y=\sum_{i=1}^{n}t_{i}x_{i}$, and here we assume the number of nonzero elements in $t_{1},...,t_{n}$ is minimized. Removing those zero coefficients in $t_{1},...,t_{n}$, we can write $y=\sum_{i=1}^{k}t_{j_{i}}x_{j_{i}}$, where the index set $\{j_{1},...j_{k}\}$ is a subset of $\{1,...,n\}$, $t_{j_{1}},...,t_{j_{k}}>0$ and $\sum_{i=1}^{k}t_{j_{i}}=1$. We only need to show that $k>p$ implies that $y$ must be an interior point of the convex hull of the point set $\{x_{1},...,x_{n}\}$.
We can write
$y-x_{j_{1}}=t_{j_{2}}(x_{j_{2}}-x_{j_{1}})+...+t_{j_{k}}(x_{j_{k}}-x_{j_{1}})=:t_{j_{2}}w_{2}+...t_{j_{k}}w_{k},$
where $t_{j_{2}},...,t_{j_{k}}>0$, $\sum_{i=2}^{k}t_{j_{i}}<1$ and $w_{i}=x_{j_{i}}-x_{j_{1}}$ for $2\leq i\leq k$. We first show that $w_{2},...,w_{k}$ must be linearly independent. Suppose $w_{2},...,w_{k}$ are not linearly independent. Then, there exist $\gamma_{2},...,\gamma_{k}\in\mathbb{R}$ such that $\sum_{i=2}^{k}\gamma_{i}w_{i}=0$ and at least one element in $\{\gamma_{2},...,\gamma_{k}\}$ is nonzero. We consider $\tilde{t}_{j_{i}}=t_{j_{i}}+\delta\gamma_{i}$. Then, $y-x_{j_{1}}=\sum_{i=2}^{k}\tilde{t}_{j_{i}}w_{i}$ and $\sum_{i=2}^{k}\tilde{t}_{j_{i}}<1$ hold for all $\delta\in\mathbb{R}$. By continuity, we can choose $\delta$ such that one element in $\{\tilde{t}_{j_{2}},...\tilde{t}_{j_{k}}\}$ is zero and all other elements are nonnegative. Then
$y-x_{j_{1}}=\tilde{t}_{j_{2}}(x_{j_{2}}-x_{j_{1}})+...+\tilde{t}_{j_{k}}(x_{j_{k}}-x_{j_{1}}),$
which contradicts the minimal choice of $k$. Therefore, we finish the proof that $w_{2},...,w_{k}$ must be linearly independent.
Since $k>p$ and $w_{2},...,w_{k}$ are linearly independent, it is only possible that $k=p+1$. For any $v$ in the unit closed ball in $\mathbb{R}^{p}$, $v$ can be written as a unique linear combination of $w_{2},...,w_{k}$. Due to the compactness of the unit closed ball in $\mathbb{R}^{p}$, there exists $M>0$ such that these absolute values of the linear combination coefficients are less than $M$. Now, we choose
$r=\frac{1}{M}\min\left\{t_{j_{2}},...t_{j_{k}},\frac{1-\sum_{i=2}^{k}t_{j_{i}}}{k-1}\right\}.$
Then, for any $y^{\prime}\in\mathbb{R}^{p}$ satisfying $\|y^{\prime}-y\|_{2}<r$, we can write
$y^{\prime}-y=\lambda_{2}w_{2}+...+\lambda_{k}w_{k},$
where $|\lambda_{i}|<rM$, and thus
$y^{\prime}-x_{j_{1}}=(t_{j_{2}}+\lambda_{2})w_{2}+...+(t_{j_{k}}+\lambda_{k})w_{k}$
where $t_{j_{2}}+\lambda_{2},...,t_{j_{k}}+\lambda_{k}>0$ and $\sum_{i=2}^{k}(t_{j_{i}}+\lambda_{i})<1$. This implies that $y$ is an interior point of the convex hull of the point set $\{x_{1},...,x_{n}\}$. ∎

### v2

@@ 5.1 Simulation under the random graph models
We first conduct a simulation study under the random graph models to evaluate the numerical performance of different fairness criteria. We consider a binary protected feature that indicates whether one is white or non-white. Specifically, we fix 80 white incompatible pairs (80%) and 20 non-white incompatible pairs (20%). Among the 80 white pairs, 56 patients (70%) are lowly-sensitized, 16 patients (20%) are moderately-sensitized, and 8 patients (10%) are highly-sensitized; and among the 20 non-white pairs, 14 patients (70%) are lowly-sensitized, 4 patients (20%) are moderately-sensitized, and 2 patients (10%) are highly-sensitized. The PRA scores are set to be 0.9, 0.45, and 0.05 for highly-sensitized, moderately-sensitized, and lowly-sensitized patients, respectively. The distribution follows the analysis in Saidman et al. (2006).
@@
We compare three fairness criteria, group fairness, individual fairness, and our newly proposed fairness criterion, using the outcome with no fairness constraints as a baseline. For our new fairness criterion, we evaluate the two parameter choices introduced in Section 3.1. For group fairness, we consider two parameter-selection methods. The first selects the largest feasible value of $\alpha$ for which (2) remains solvable, following Dickerson et al. (2014). This corresponds to imposing the strongest possible constraint, requiring us to maximize the number of matched highly sensitized patients. The second method selects the largest $\alpha$ that preserves the overall utility achieved in (1), as in Freedman et al. (2020). In other words, among all exchange plans that maximize the objective in (2), we choose the one that yields the greatest number of matched highly sensitized patients. For individual fairness, we adopt two analogous parameter-selection approaches. Specifically, we minimize the $L_{1}$ norm of the difference vector $\pi-\overline{\pi}$ while retaining at least 80% and 100% of the maximum achievable utility, respectively. These correspond to a stronger and a weaker fairness constraint.
Figure 2 presents the average selection rates of different fairness criteria within each subgroup. From Figure 2, group fairness works to increase the selection rates of highly-sensitized patients, and individual fairness works to balance the selection rates of the six subgroups. Differently, our fairness balances the selection rates of the white and non-white patients within each sensitization stratum, instead of the selection rates of all six subgroups. Our stronger fairness constraint can almost eliminate the absolute difference in selection rates, while our weaker fairness constraint can also significantly reduce the absolute difference in selection rates compared to other methods. For the price of fairness, the relative utility loss is around 6.3% for our stronger fairness constraint and 0.7% for our weaker fairness constraint.
@@ 5.2 UNOS data analysis
Figure 3 reports the average selection rates within each subgroup. From Figure 3, group fairness, individual fairness, and our new fairness criterion all enhance equitable access to transplant opportunities according to their specific fairness criteria.
@@ 5.3 Experiment on selection probability prediction
We next conduct a numerical experiment to evaluate the prediction accuracy of the selection probability. We assume there is an underlying population consisting of 80% white and 20% non-white incompatible pairs, following the same blood type distributions and sensitization distributions in Section 5.1. We assume the historical data $\{\tilde{v}_{1},\cdots,\tilde{v}_{N_{0}}\}$, current data $\{v_{1},\cdots,v_{N_{1}}\}$, and future data are independently sampled from the population. We fix $N_{0}=200$, $N=100$, $B=1000$, and vary $N_{1}$ in $\{20,40,60,80\}$. For illustration purposes, we focus on our new fairness criterion.
Table 1 presents the numerical results of the selection probability prediction. As $L$ increases, the prediction accuracy improves, evidenced by a lower mean squared error (MSE) and a coverage rate of the prediction interval approaching 95%. This improvement occurs because the size of the unobserved future data, $N-N_{1}$, decreases, which makes the prediction easier. Although these results indicate the method’s validity, there is a bias in the prediction interval due to the finite size of $N_{0}$. When $N_{0}$ is large, the set $\{\tilde{v}_{1},\cdots,\tilde{v}_{N_{0}}\}$ closely approximates the underlying population distribution; while when $N_{0}$ is small, the distribution of $\{\tilde{v}_{1},\cdots,\tilde{v}_{N_{0}}\}$ may differ from the underlying population distribution.
@@
MSE 0.079 0.056 0.045 0.038
Coverage 0.977 0.965 0.962 0.958
Width 0.445 0.332 0.265 0.227
Table 1: Accuracy of the predicted selection probability of $v_{1},\cdots,v_{N_{1}}$. The results are averaged over 100 data replications.
@@ Appendix A Proofs
Fix $k\geq 2$ and $p\in(0,1)$. For each $n$, let $(A_{1},\dots,A_{k})$ be a $k$-partite undirected random graph with $|A_{i}|=n$ for all $i$, where every potential edge between distinct parts is present independently with probability at least $p$.
Then, with probability 1o(1)1-o(1) as $n\to\infty$, the graph contains $n$ vertex-disjoint copies of the complete $k$-partite graph with one vertex in each part; that is, there exist $n$ disjoint $k$-tuples, each consisting of one vertex from every part $A_{i}$, such that all $\binom{k}{2}$ inter-part edges among the vertices in each $k$-tuple are present.
@@
We build the statement in four steps.
Step 1: Bipartite building block. Let $X,Y$ be disjoint sets of size $n$, and consider a bipartite graph in which each edge $(x,y)$ is present independently with probability at least $q>0$. By monotone coupling with $G(n,n,q)$, this random graph stochastically dominates an Erdős–Rényi bipartite graph. It is classical that $G(n,n,q)$ has a perfect matching with probability 1o(1)1-o(1) as $n\to\infty$, and “having a perfect matching” is monotone in the edge set. Hence the original bipartite graph also has a perfect matching with probability 1o(1)1-o(1).
Step 2: Uniform representation and filtration. For each unordered pair $(u,v)$ with $u\in A_{i}$, $v\in A_{j}$, $i\neq j$, let $U_{(u,v)}$ be independent $\mathrm{Unif}[0,1]$. If the designated probability of the edge $(u,v)$ is $p_{(u,v)}\in[p,1]$, we declare $(u,v)$ present iff $U_{(u,v)}\leq p_{(u,v)}$. This yields exactly the model in the lemma.
For $t=1,\dots,k-1$, define
$\mathcal{F}_{t}:=\sigma\bigl(U_{(u,v)}:u\in A_{i},\ v\in A_{j},\ i\neq j,\ \{u,v\}\subseteq A_{1}\cup\cdots\cup A_{t+1}\bigr),$
which reveals all edges whose endpoints lie in $A_{1},\dots,A_{t+1}$.
Step 3: Inductive construction of $(t+1)$-tuples. We build the desired $k$-tuples in k1k-1 rounds. At the end of round $t\geq 2$ we will have $n$ disjoint sets $B^{(t)}_{1},\dots,B^{(t)}_{n}$, each containing one vertex from each of $A_{1},\dots,A_{t}$ and inducing a complete $t$-partite subgraph.
Round $t=2$. Consider the bipartite graph between $A_{1}$ and $A_{2}$. All edges are independent and present with probability at least $p$, so by Step 1 it has a perfect matching with probability 1o(1)1-o(1). Fix such a matching and write its edges as $(x_{j},y_{j})$, $j=1,\dots,n$, with $x_{j}\in A_{1}$, $y_{j}\in A_{2}$, and define
$B^{(2)}_{j}:=\{x_{j},y_{j}\},\qquad B_{2}:=\{B^{(2)}_{1},\dots,B^{(2)}_{n}\}$.
Inductive step. Assume for some $t\in\{2,\dots,k-1\}$ that we have disjoint sets
$B_{t}=\{B^{(t)}_{1},\dots,B^{(t)}_{n}\}$
such that each $B^{(t)}_{j}$ contains exactly one vertex from each of $A_{1},\dots,A_{t}$ and induces a complete $t$-partite subgraph. Then $B_{t}$ is $\mathcal{F}_{t-1}$-measurable.
Define a bipartite graph $\Gamma_{t}$ whose left part is $B_{t}$ and whose right part is $A_{t+1}$. For $B\in B_{t}$ and $a\in A_{t+1}$, put an edge between $B$ and $a$ in $\Gamma_{t}$ if and only if $a$ is adjacent to every vertex in $B$ in the original graph, i.e.
$(B,a)\in E(\Gamma_{t})\quad\Longleftrightarrow\quad U_{(v,a)}\leq p_{(v,a)}\ \text{for all }v\in B.$
Conditional on $\mathcal{F}_{t-1}$, $B_{t}$ is fixed and the variables $\{U_{(v,a)}:v\in B\}$ are independent, with $|B|=t$ and $p_{(v,a)}\geq p$. Hence
$\mathrm{Pr}\bigl((B,a)\in E(\Gamma_{t})\mid\mathcal{F}_{t-1}\bigr)=\prod_{v\in B}p_{(v,a)}\geq p^{t}=:q_{t}>0.$
For distinct pairs $(B,a)\neq(B^{\prime},a^{\prime})$, the sets of uniforms $\{U_{(v,a)}:v\in B\}$ and $\{U_{(v^{\prime},a^{\prime})}:v^{\prime}\in B^{\prime}\}$ are disjoint, so the edge indicators in $\Gamma_{t}$ are independent, each with conditional probability at least $q_{t}$.
Thus, conditional on $\mathcal{F}_{t-1}$, $\Gamma_{t}$ satisfies the assumptions of Step 1 with parameter $q_{t}$, and since $|B_{t}|=|A_{t+1}|=n$, it has a perfect matching with probability 1o(1)1-o(1). On that event, write the matching as $(B^{(t)}_{j},a_{j})$, $j=1,\dots,n$, and set
$B^{(t+1)}_{j}:=B^{(t)}_{j}\cup\{a_{j}\},\qquad B_{t+1}:=\{B^{(t+1)}_{1},\dots,B^{(t+1)}_{n}\}$.
Each $B^{(t+1)}_{j}$ is then a complete $(t+1)$-partite subgraph with one vertex in each of $A_{1},\dots,A_{t+1}$.
Step 4: Union bound over rounds. There are k1k-1 rounds. For each fixed $t$, the conditional failure probability is $o(1)$ as $n\to\infty$, hence the unconditional failure probability is $o(1)$. Since $k$ is fixed, a union bound shows that with probability 1o(1)1-o(1) all rounds succeed, yielding $n$ disjoint $k$-tuples as in the statement. ∎